OpenStudy (anonymous):

Suppose that the functions g and h are defined for all real numbers x as follows. g(x)-2x^2 and h(x)=4x write the expressions for (g+h)(x) and (g-h)(x) and evaluate (g*h)(2)

OpenStudy (mathmale):

What you're attempting here is called "addition of functions." (g+h)(x) is the sum of g(x) = 2x^2 and h(x)=4x. This sum is found by adding like terms: g(x) = 2x^2 +h(x) = + 4x ------------------ g(x)+h(x) = 2x^2 + 4. That's it. Try finding (g-h)(x). I'll be back, but it may be 15-30 minutes or more. Good luck!

OpenStudy (isaiah.feynman):